A body of mass12kg travelling at 42m/s collides with a body of mass 18kg at rest calculate their common velocity if the two bodies coalesce after coalition
1
Expert's answer
2020-10-21T09:53:26-0400

We can find the common velocity of two bodies after the collision from the Law of Conservation of Momentum:


"m_1v_1+m_2v_2=(m_1+m_2)v_{c},"

here, "v_{c}" is the common velocity of two bodies after the collision, "m_1=12\ kg" is the mass of the first body, "v_1=42\ ms^{-1}"is the velocity of the first body, "m_2=18\ kg" is the mass of the second body, "v_2=0\ ms^{-1}"is the velocity of the second body.

Then, from this equation we can find the common velocity of two bodies after the collision:


"v_c=\dfrac{m_1v_1+m_2v_2}{(m_1+m_2)},""v_c=\dfrac{12\ kg\cdot 42\ ms^{-1}+18\ kg\cdot 0\ ms^{-1}}{(12\ kg+18\ kg)}=16.8\ ms^{-1}."

Answer:

"v_c=16.8\ ms^{-1}."